CDB stock is currently priced at $63. The company will pay a dividend of $4.81 next year and investors require a return of 11.1 percent on similar stocks. What is the dividend growth rate on this stock?
Answer #1

Ans 3.47%

P0 = Price of Share
D1 = Current Dividend
Ke = Cost of Equity
g = growth rate
P0 = D1 / (Ke - g)
63 = 4.81 / (11.1%- g)
11.1% - g = 4.81 / 63
11.1% - g = 7.63%
g = 11.1% - 7.63%
g = 3.47%